### v2

# The $\infty$-S test via quantile affine LASSO

Abstract
The nonparametric sign test dates back to the early 18th century with a data analysis by John Arbuthnot. It is an alternative to Gosset’s more recent $t$-test for consistent differences between two sets of observations. Fisher’s $F$-test is a generalization of the $t$-test to linear regression and linear null hypotheses. Only the sign test is robust to non-Gaussianity. Gutenbrunner et al. (1993) derived a version of the sign test for linear null hypotheses in the spirit of the F-test, which requires the difficult estimation of the sparsity function. We propose instead a new sign test called $\infty$-S test via the convex analysis of a point estimator that thresholds the estimate towards the null hypothesis of the test.
Keywords: LASSO, nonparametric statistical test, rescaling, robustness, total variation.
1 Introduction
Given a goodness-of-fit measure such as the log-likelihood, the score test calculates the $\ell_{2}$-norm (weighted by the inverse of Fisher information matrix) of the gradient evaluated at the null hypothesis $H_{0}$ and evaluates whether it is abnormally large compared to a chi-squared distribution. The idea is that if the restricted (under the constraint defined by $H_{0}$) estimator is close the free optimum of the goodness-of-fit measure, the score should not differ from zero significantly. The main advantage of the score test is that it requires only the calculation of the restricted estimate. The score test is equivalent to the Lagrange multiplier test (Silvey, 1959), and to the F-test (Kohler, 1982) for Gaussian linear regression which assumes the data $(X,\boldsymbol{y})$ are related according to
$\boldsymbol{y}=X\boldsymbol{\beta}+\boldsymbol{e},$ (1)
where $X$ is an $n\times p$ matrix and $\boldsymbol{y}\in{\mathbb{R}}^{n}$ is a response, $\boldsymbol{e}$ are errors i.i.d. sampled from a Gaussian distribution ${\rm N}(0,\sigma^{2})$. Given an $m\times p$ matrix $A$ of rank $m$ and a vector $\boldsymbol{b}\in{\mathbb{R}}^{m}$, the F-test considers the linear null and alternative hypotheses
$H_{0}:\ A\boldsymbol{\beta}=\boldsymbol{b}\quad{\rm against}\quad H_{1}:\ A\boldsymbol{\beta}\neq\boldsymbol{b}.$ (2)
We assume that ${\rm rank}(X)=\min(n,p)$ and ${\rm rank}(X)-m<n$, for instance, the situations where $p<n$ or where $p\geq n$ and $m=p$. To derive a pivotal statistic that depends not only on $(X,\boldsymbol{y})$ but also on $(A,\boldsymbol{b})$, Fisher’s F-test uses the Lagrange multiplier technique to define the statistic
$F=\frac{(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})^{\rm T}[A(X^{\rm T}X)^{-1}A^{\rm T}]^{-1}(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})/m}{(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})^{\rm T}(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})/(n-p)},$
where $\hat{\boldsymbol{\beta}}=(X^{\rm T}X)^{-1}X^{\rm T}\boldsymbol{Y}$ is the least squares solution of the full model. Under the null $H_{0}$ and for Gaussian noise, the statistic is pivotal with distribution $F_{m,n-p}$. The F-test is not robust to non-Gaussian errors however. Consequently the level of the F-test may not match the nominal level.
Assuming a noise distribution through the choice of a likelihood, the tests discussed above are parametric and may therefore be sensitive to deviation from the assumed distribution. Another class of tests is nonparametric. The oldest nonparametric test (beginning of 18th century) is certainly the sign test employed in the same setting as the paired $t$-test. The advantage of the sign test is that it does not assume Gaussianity, and therefore is robust to deviation from this assumption. Nonparametric tests often make the relaxed assumption that the errors $\boldsymbol{e}$ are i.i.d. sampled from a random variable with zero median. The advantage of nonparametric tests is that they match the nominal level under weak assumptions. The drawback is their lack of power; the signed-rank test is a weighted version of the sign test that gives more power to it, while remaining nonparametric (Wilcoxon, 1945).
Gutenbrunner et al. (1993) extended the sign-based tests for testing equal median of several populations to the linear regression setting (1), where now the errors $\boldsymbol{e}$ are sampled from a random variable with zero median and scale parameter $\xi$ (e.g., the standard deviation for the Gaussian distribution and the median absolute deviation for the Cauchy distribution). Splitting $X=[X_{1},X_{2}]$ and correspondingly $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ with $X_{1}$ and $X_{2}$ have $n$ rows and $p_{1}$ and $p_{2}$ columns, Gutenbrunner et al. (1993) allow to test the exclusion null and alternative hypotheses
$H_{0}:\ \boldsymbol{\beta}_{2}=\boldsymbol{0}\quad{\rm against}\quad H_{1}:\ \boldsymbol{\beta}_{2}\neq\boldsymbol{0}.$ (3)
Letting $O_{d_{1},d_{2}}$ be the zero matrix of dimension $d_{1}\times d_{2}$ and $I_{d_{1}}$ be the identity matrix of dimension $d_{1}\times d_{1}$, these exclusion hypotheses correspond to the general hypotheses (2) with $A=[O_{p_{2}\times p_{1}},I_{p_{2}}]$, $\boldsymbol{b}=\boldsymbol{0}_{p_{2}}$ and $m=p_{2}$. In the spirit of the F-test, they derive a statistic
$T_{n}=S_{n}^{\rm T}Q_{n}^{-1}S_{n}/C_{\varphi}\rightarrow_{d}\chi^{2}_{p_{2}},$ (4)
where $S_{n}$ is the score under $H_{0}$ and $C_{\varphi}$ is some constant; see Gutenbrunner et al. (1993, eq. (2.9)) for details. Their test statistic is the weighted $\ell_{2}$-norm of the score, and the rescaling by $Q_{n}^{-1}$ makes the statistic asymptotically pivotal and $\chi^{2}$ distributed. For that reason we call their test the $\chi^{2}$ sign test. In practice $Q_{n}$ must be estimated which entails estimating the sparsity function $s(\tau)=(f(F^{-1}(\tau)))^{-1}$, where $f$ and $F$ are the unknown density and c.d.f. of the errors, respectively, and $\tau\in(0,1)$ is the quantile of interest; this additional nuisance parameter prevents the test statistic to be close to the asymptotic $\chi^{2}$ distribution even for large sample size $n$ and makes the implementation unstable, as we observe in our empirical experiments in Section 5.
Interestingly the exclusion hypotheses (3) allow to test the general hypotheses (2) for any $(A,\boldsymbol{b})$ using the following transformation. Letting $\tilde{A}$ be a $(p-m)\times p$ matrix that completes $A$ such that $[\tilde{A}^{\rm T},A^{\rm T}]$ is nonsingular and $A\tilde{A}^{\rm T}=O_{m\times(p-m)}$, they transform the data to $\tilde{\boldsymbol{y}}=\boldsymbol{y}-XA^{\rm T}(AA^{\rm T})^{-1}\boldsymbol{b}$, $\tilde{X}_{1}=X\tilde{A}^{\rm T}(\tilde{A}\tilde{A}^{\rm T})^{-1}$ and $\tilde{X}_{2}=XA^{\rm T}(AA^{\rm T})^{-1}$ and apply their test with $(\tilde{X}_{1},\tilde{X}_{2},\tilde{\boldsymbol{y}})$ in place of $(X_{1},X_{2},\boldsymbol{y})$ to test (2). Mathematically elegant, this transformation is not innocuous because computationally cumbersome. Indeed this transformation requires to find a matrix $\tilde{A}$ with the desired properties and invert $\tilde{A}\tilde{A}^{\rm T}$. Take the example where one wants to test $\beta_{1}=\beta_{2}$: $A$ is a $1\times p$ matrix, so $\tilde{A}$ is $(p-1)\times p$ and $\tilde{A}\tilde{A}^{\rm T}$ is $(p-1)\times(p-1)$, which may be difficult to invert with today’s data where $n$ and $p$ are large. We see that this indirect way for handling the general linear hypotheses is tedious.
Like the F-test, we derive instead a statistic that directly tackles the desired null hypotheses (2), but as opposed to the F-test, we do not seek an asymptotic $\chi^{2}$ distribution. Our derivation of the $\infty$-S test is based on the least absolute deviation version of the affine LASSO point estimator (Sardy et al., 2022). Affine LASSO is a regularized estimator which regularization parameter can be seen as a Lagrange multiplier to constrain the null hypothesis. So the $\infty$-S test is both a score test and a Lagrange multiplier test.
The article is organized as follows. In Section 2, we derive the $\infty$-S test by defining the affine LAD LASSO (least absolute deviation least absolute shrinkage and selection operator), deriving its zero-thresholding function, and defining the sign-based test statistic that is asymptotically pivotal with respect the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO. In Section 4, we extend the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile affine LASSO estimator. In Section 5, we apply the $\infty$-S test to four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
2 The $\infty$-S test
The affine LAD-LASSO estimator is a solution to
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}.$ (5)
The larger $\lambda$, the more this estimator thresholds towards zero the entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ corresponding to the null hypothesis. As stated in the following theorem, there is an interval $[\lambda_{0},\infty)$ for which the solution to (5) satisfies the condition of $H_{0}$ for any $\lambda\in[\lambda_{0},\infty)$. The smallest such $\lambda$ is a function of the data and has a closed form expression called the zero-thresholding function.
Theorem 1.
Let
$\hat{\boldsymbol{\beta}}_{H_{0}}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\quad{\rm s.t.}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (6)
The zero-thresholding function of the affine LAD-LASSO primal problem (5) is
$\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty},$ (7)
where ${\boldsymbol{\omega}}$ is the dual variable associated to (5), which are the sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}$ of the constrained LAD (6).
Since satisfying the conditions of $H_{0}$ are equivalent to $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$ for any norm, the zero-thresholding function (7) is the (scalar) Lagrange multiplier to satisfy the constraint $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$; using the $\ell_{1}$-norm, the Lagrange multiplier that achieves this is finite. Instead of the $\ell_{1}$-norm, we could have used the $\ell_{2}$-norm corresponding to affine LAD-group LASSO (Yuan and Lin, 2006); the dual of the $\ell_{2}$-norm being the $\ell_{2}$-norm, the corresponding zero-thresholding function would have been $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{2}$. More generally, one can use the $\ell_{q}$-norm, which corresponding zero-thresholding function is $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{q/(q-1)}$.
The zero-thresholding function leads to the statistic of the $\infty$-S test, which is asymptotically pivotal, as stated in the following theorem.
Theorem 2.
Assuming the LAD coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$, where $\boldsymbol{Y}$ is the response random vector under $H_{0}$ is asymptotically pivotal.
The test statistic $S$ involves the score of the LAD loss under $H_{0}$ like the statistic $T_{n}$ of Gutenbrunner et al. (1993) with two noticable differences: the test statistic $S$ uses the infinite norm instead of the weighted $\ell_{2}$-norm, and its distribution does not require estimating the sparsity function of Gutenbrunner et al. (1993). The sparisty function is only needed to make the statistic asymptotically $\chi^{2}$under $H_{0}$, not to make the statistic pivotal. Indeed, as Theorem 2 shows, the score is already asymptotically pivotal. There is an important change of paradigm between the generalization of the sign test by Gutenbrunner et al. (1993) and our new $\infty$-S test: their test has an asymptotic listed distribution (e.g., $F$ or $\chi^{2}$), while our test has an unlisted distribution. Nevertheless, with today’s computer, one can easily sample the distribution of the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ by Monte Carlo and estimate the test critical value as the empirical $(1-\alpha)$-quantile of $S$. The p-value is also estimated as the percentage of Monte Carlo simulated random variables larger than $\lambda_{0}(\boldsymbol{y},X,A)$.
The asymptotic Gaussianity of the LAD coefficient estimates required by Theorem 2 has been proved by Pollard (1991), so we can define the $\infty$-S test.
Definition 1 (The $\infty$-S test).
Consider the linear model (1). To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}=F_{S}^{-1}(1-\alpha)$, where $S=\lambda_{0}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}$ is given in (7). The $\infty$-S test function is defined by
$\phi({\bf y},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}(\boldsymbol{y},X,A)\leq c_{\alpha}\\ 1&{\rm otherwise}\end{array}\right.,$
where ${\boldsymbol{y}}$ is the data response.
Based on Theorem 2, the level of the $\infty$-S test matches asymptotically the nominal level $\alpha$. The $\infty$-S test looks at the sign of the residuals, not their amplitude, causing some lack of power of the $\infty$-S test with respect to the F-test for Gaussian models. Weighting the sign $\boldsymbol{\omega}$ of the residuals by the rank of the absolute value of the respective residuals leads to the $\infty$-rankS test that has more power.
To completely define the test, the rescaling of $(A,\boldsymbol{b})$ must be performed to achieve homogeneous power, that it, to not favor certain alternative hypotheses to the detriment of others. Indeed, the matrices $A$ and $X$ create heteroschedasticity in the random vector $\boldsymbol{W}$ in the test statistics $S=\|\boldsymbol{W}\|_{\infty}$ with $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ before taking the maximum of all entries. Consequently, under $H_{1}$, the entries of $\boldsymbol{W}$ with small deviations will be hidden behind the entries with large deviations. To be democratic between potential entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ not being equal to zero, we propose the following rescaling.
Definition 2 (Homopower rescaling).
Consider the $\infty$-S test with test statistic $S=\|\boldsymbol{W}\|_{\infty}$, where $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ and ${\boldsymbol{\omega}}$ are the sign of the residuals of the constrained LAD under $H_{0}$. Let $F_{W_{k}}$ be the marginal cumulative distribution function of the $k$th entry of $\boldsymbol{W}$ for $k=1,\ldots,m$. Letting $D_{(A,X,{\bf b},\alpha)}$ be the diagonal matrix with $k$th diagonal entry equal to $d_{k}:=F_{W_{k}}^{-1}(1-\alpha)$, homopower rescaling rescales $A$ to $D_{(X,A,{\bf b},\alpha)}A$ and $\boldsymbol{b}$ to $D_{(X,A,{\bf b},\alpha)}\boldsymbol{b}$ before applying the test.
3 Optimization
The cost function of the LAD optimization
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}$ (8)
is convex, and it is well known that the optimization can be transformed into a linear program (Koenker and Bassett, 1978). The $\infty$-S test requires solving the constrained LAD optimization (6) that can be transformed into an unconstrained LAD optimization of the form (8).
Theorem 3.
Solving (6) is equivalent to solving (8) with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$, where $A_{1}$ are $m$ linearly independent columns of $A=:[A_{1},A_{2}]$ and $X=:[X_{1},X_{2}]$ is the corresponding decomposition.
The proof is straightforward: choose $m$ linearly independent columns of $A$, call them $A_{1}$ and the remaining ones $A_{2}$, so that $A=[A_{1},A_{2}]$, permuting $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ accordingly. Hence the linear constraints leads to $\boldsymbol{\beta}_{1}=A_{1}^{-1}(\boldsymbol{b}-A_{2}\boldsymbol{\beta}_{2})$ and $\boldsymbol{y}-X\boldsymbol{\beta}$ becomes $\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{2}$ with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$.
The affine LAD-LASSO optimization (5) can also be transformed into an unpenalized LAD of the form (3) since it is equivalent to solving
$\min_{{\boldsymbol{\beta}}\in{\mathbb{R}}^{p}}\|\left(\begin{array}[]{c}{\boldsymbol{y}}\\ {\lambda{\boldsymbol{b}}}\end{array}\right)-\left(\begin{array}[]{c}X\\ {\lambda A}\end{array}\right){\boldsymbol{\beta}}\|_{1}.$
Another way to solve the constrained LAD optimization (6) is to solve the affine LAD LASSO (5) for $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ of Theorem 1 to guarantee complete thresholding, that is, forcing the linear constraint; this is akin to the exact penalty method of Di Pillo and Grippo (1989). The following theorem links the dual of the affine LAD LASSO optimization (5) and the constrained LAD optimization.
Theorem 4.
The dual variable $\boldsymbol{\omega}$ of Theorem 1 corresponding to the primal affine LAD LASSO optimization (5) is the dual variable of the LAD optimization (8) with $\boldsymbol{y}_{-A}$ and $X_{-A}$ of Theorem 3.
So the $\infty$-S test can straightforwardly be implemented with current softwares, in particular using the quantreg package (Koenker, 2024) in R.
4 The quantile $\infty$-Sτ test
Given $\tau\in(0,1)$, consider the $\tau$-conditional quantile vector ${\boldsymbol{q}}^{\tau}$ of a response vector ${\boldsymbol{Y}}\mid X$, where $X$ is the regression matrix. The linear model assumes
${\boldsymbol{q}}^{\tau}=X{\boldsymbol{\beta}}.$ (9)
To estimate the parameters ${\boldsymbol{\beta}}$ based on a sample ${\boldsymbol{y}}$ from ${\boldsymbol{Y}}\mid X$, quantile regression (Koenker and Bassett, 1978) considers solving
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{\rho_{\tau}},$ (10)
where the residuals are $\boldsymbol{r}=\boldsymbol{y}-X\boldsymbol{\beta}$ and $\|\boldsymbol{r}\|_{\rho_{\tau}}=\sum_{i=1}^{n}\rho_{\tau}(r_{i})$ with
$\rho_{\tau}(r_{i})=\left\{\begin{array}[]{rl}(\tau-1)r_{i}&r_{i}<0\\ \tau r_{i}&r_{i}\geq 0\end{array}\right.$ (11)
is the tilted $\ell_{1}$ loss. In particular $\|\boldsymbol{r}\|_{\rho_{1/2}}=\|{\boldsymbol{r}}\|_{1}/2$.
To test the null hypothesis against the alternative hypothesis (2), in the spirit of quantile regression, one derives the quantile $\infty$-S test to not necessarily fit the data to the median (that is $\tau=1/2$), but to any $\tau$-quantile for some fixed $\tau\in(0,1)$. To that aim, we define affine $\rho_{\tau}$-LASSO as a solution to
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}$ (12)
for some $\lambda>0$ and $\tau\in(0,1)$. This estimator has a zero-thresholding function given by the following theorem.
Theorem 5.
Given $\tau\in(0,1)$, let
$\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})\quad{\rm s.t.}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (13)
The zero-thresholding function of the affine $\rho_{\tau}$-LASSO primal problem (12) is
$\lambda_{0}^{\tau}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty},$ (14)
where ${\boldsymbol{\omega}}^{\tau}$ is the dual variable associated to (12), which are the weighted sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{y}}$ of the constrained minimum $\rho_{\tau}$ (13).
The zero-thresholding function leads to the statistic of the $\infty$-Sτ test, which is asymptotically pivotal, as stated in the following theorem.
Theorem 6.
Given $\tau\in(0,1)$, assuming the quantile regression coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$, where $\boldsymbol{Y}$ is the response random vector under $H_{0}$ is asymptotically pivotal.
Asymptotic Gaussianity has been proved by Koenker and Bassett (1978), so the $\infty$-Sτ test can now be defined.
Definition 3 (The $\infty$-Sτ test).
Given $\tau\in(0,1)$, consider the linear model (9) for the quantiles. To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}^{\tau}=F_{S^{\tau}}^{-1}(1-\alpha)$, where $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}^{\tau}$ is given in (14). The quantile $\infty$-Sτ test function is defined by
$\phi({\bf y},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}^{\tau}(\boldsymbol{y},X,A)\leq c_{\alpha}^{\tau}\\ 1&{\rm otherwise}\end{array}\right.,$
where ${\boldsymbol{y}}$ is the data response.
Based on Theorem 6, the level of the $\infty$-Sτ test is asymptotically the nominal level $\alpha$. Weighting the sign $\boldsymbol{\omega}^{\tau}$ of the residuals by the ranks of the weighted residuals $|\boldsymbol{r}|\cdot\rho_{\tau}(\boldsymbol{r})$ leads to the $\infty$-rankSτ test.
5 Application
Our simulations and application are done with the R software and rely on the quantreg package (Koenker, 2024). We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ necessary to calculate the zero-thresholding function. The critical values are estimated by Monte Carlo simulation with $10^{4}$ runs in all our applications of the $\infty$-S test. The levels and power curves of the tests considered are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the $\chi^{2}$ sign test whenever it returned a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
5.1 The paired $\infty$-S test
Consider two paired measurements $\{(u_{i},v_{i})\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu_{i}+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu_{i}+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
where the errors $\tilde{\epsilon}$ and $\breve{\epsilon}$ are assumed independent with a median equal to zero. To test
$H_{0}:\delta=0\quad{\rm against}\quad H_{1}:\delta\neq 0$
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{v}}-{\boldsymbol{u}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. Then the test-statistic value is $s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{v}}-{\boldsymbol{u}})|=|\sum_{i=1}^{n}(1(v_{i}>u_{i})-1(v_{i}<u_{i}))|$. Recall that the sign test’s statistic value is $\tilde{s}=\sum_{i=1}^{n}1(v_{i}>u_{i})$. But $\sum_{i=1}^{n}1(v_{i}>u_{i})+\sum_{i=1}^{n}1(v_{i}<u_{i})=n$, so $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the existing sign test.
5.2 The unpaired $\infty$-S test
Consider two unpaired measurements $\{u_{i}\}_{i=1,\ldots,n}$ and $\{v_{i}\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
where the errors $\tilde{\epsilon}$ and $\breve{\epsilon}$ are assumed independent with a median equal to zero. To test
$H_{0}:\delta=0\quad{\rm against}\quad H_{1}:\delta\neq 0$
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X{\boldsymbol{\beta}}+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}^{\rm T}=({\boldsymbol{u}}^{\rm T},{\boldsymbol{v}}^{\rm T})$,
$X=\left(\begin{array}[]{cc}\begin{array}[]{c}{\boldsymbol{1}}_{n}\\ {\boldsymbol{1}}_{n}\end{array}&\begin{array}[]{c}{\boldsymbol{0}}_{n}\\ {\boldsymbol{1}}_{n}\end{array}\end{array}\right),$
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is $s=\max(0,|\sum_{i=1}^{n}1(v_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(v_{i}<m_{\boldsymbol{y}})|)$. This is the sign test for unpaired data, also known as the median test (Brown and Mood, 1951).
Figure 1: Power of the unpaired $\infty$-Sτ test (same as median test) and $\infty$-rankSτ test for $\tau=0.5$, the F-test and asymptotically $\chi^{2}$ sign test as a function of the shift $\delta\in[0,2]$ between the two populations. The horizontal dotted line is the nominal level $\alpha=0.05$. Sample size $n=20$ (left plot) and $n=100$ (right plot).
We compare the level and the power of four tests ($\infty$-S or median test, $\infty$-rankS test, F-test and $\chi^{2}$ sign test) on two Monte Carlo simulations with $n=20$ and $n=100$ and with Student errors with 3 degrees of freedom. Figure 1 shows that all tests satisfy the nominal level of $\alpha=0.05$ and the F-test looks quite robust to non-Gaussian errors although losing some power. The $\chi^{2}$ sign test seems to overshoot the level, even for $n$ large, however.
5.3 Homopower rescaling
We perform a toy simulation to outline the importance of homopower rescaling (see Definition 2) the two components $A$ and $\boldsymbol{b}$ of the linear null hypothesis (2). Here $n=100$, $p=20$, $m=2$, $\tau=0.5$, $\boldsymbol{b}=\boldsymbol{0}_{m}$ and $A=[C,O_{m\times(p-m)}]$ with
$C=\left(\begin{array}[]{ccc}3&0\\ 0&1\end{array}\right).$
The regression matrix $X$ has entries sampled from a standard Gaussian random variable, the null model is $H_{0}:\ \boldsymbol{\beta}=\boldsymbol{0}$, and we consider two alternative hypotheses
$H_{1,1}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{1}\quad{\rm and}\quad H_{1,2}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{2},$
where $\boldsymbol{e}_{i}$ is the $i$th canonical vector. With the unrescaled $\infty$-S test, one expects the power to be low under $H_{1,1}$ and good under $H_{1,2}$, while with the rescaled $\infty$-S test, one expects the power to be high under both $H_{1,1}$ and $H_{1,2}$. Figure 2 illustrates the improvement thanks to homopower rescaling. Not only the $\infty$-S test, but many tests would profit from this rescaling in terms of power. This example is certainly a toy example, but what we observe here also occurs naturally owing to the relation between $A$ and $X$ that is not controlled by the user, except possibly in ANOVA. So homopower rescaling should be employed.
Figure 2: Homopower rescaling. Top: boxplots of Monte Carlo simulated bivariate ($m=2$) random vector $W$ unrescaled (left) and rescaled (right). Bottom: power plots of the unrescaled (red) and rescaled (black) $\infty$-S test under $H_{1,1}$ (left) and $H_{1,2}$ (right).
5.4 The total variation $\infty$-Sτtest
Total variation (Rudin et al., 1992) aims at detecting jumps in a noisy signal or a time series ${\bf y}=\boldsymbol{\beta}+\boldsymbol{e}$, which corresponds to model (1) with $X=I_{n}$. A jump occurs if $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\neq 0$. Call $A$ the $(n-1)\times n$ sparse matrix with $-1$ and $1$ entries such that $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\ =\|A\boldsymbol{\beta}\|_{1}$; $A$ is full row rank. To test for jump in the median or any $\tau$-quantile, one can apply the $\infty$-Sτ test or the $\infty$-rankSτ test at the desired $\tau$-quantile.
We apply the quantile total variation $\infty$-Sτ test to analyse the time series of the amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of $0.0132$. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic (by Monte Carlo sampling of the test statistic $S^{\tau}$ under the null hypothesis) and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Right below this plot, Figure 3 shows the quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ and the zero-thresholding function $\lambda_{0}^{\tau}(\boldsymbol{y},X,A)$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the quantile affine LASSO estimate for $\lambda=c_{\alpha}^{\tau}$ (bottom), which corresponds to the quantile universal threshold estimate (Giacoetal17) of the $\tau$-quantile. The single jump points to a potential change of regime moving to a less volatile return.
Figure 3: AMZN daily log-returm time series analysis. Top left plot: estimated density function of $f_{S^{\tau}}$ and critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$ by Monte Carlo simulation. Bottom left: regression quantile affine LASSO path for $\tau=0.1$ with critical value $c_{\alpha}^{\tau}$ and zero-thresholding function $\lambda_{0}(\boldsymbol{y},X,A)$ evaluated at $\boldsymbol{y}$. Top right plot: AMZN time series data and estimate jump locations. Bottom right: estimated $\tau$-quantiles.
5.5 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, $A$ is the first $m=5$ rows of the finite difference matrix of Section 5.4 and ${\boldsymbol{\beta}}$ are $p$ samples from the standard Gaussian distribution; $\boldsymbol{b}$ is calculated as $A{\boldsymbol{\beta}}$. In a robustness analysis of the F-test, Ali and Sharma (1996) observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom.
Figure 4: Level of the $\infty$-S test, $\infty$-rankS test and F-test as a function of the degrees of freedom of the Student errors. The dotted line is the nominal level $\alpha=0.01$.
Over-shooting the nominal level leads to too many type I errors, hence rejecting too many null hypotheses, that is, making too many false discoveries. Here the two $\infty$-S tests match the nominal level, regardless of the degree of freedom of the Student errors. For this simulation, the implementation of the quantreg package does not return a p-value for $\chi^{2}$ sign test.
6 Conclusions
The $\infty$-Sτ test allows to test general linear null hypotheses for linear models and for any quantile $\tau$. The test function is simple to implement and can handle high-dimensional problems. We emphasize the $\infty$-S test which corresponds to the $\ell_{1}$-LASSO penalty, and showed that other norms than the $\ell_{1}$-norm could be employed, leading to other tests. Sardy et al. (2022) show that using $\ell_{1}$ leads to high power under sparse alternatives (that is, most entries, but not all, of $A\boldsymbol{\beta}-\boldsymbol{b}$ are zero when $H_{1}$ is true), and, under dense alternatives (that is, most or all entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ are non-zero), that the sign test based on $\ell_{2}$ has more power than the $\infty$-S test. To get nearly the best power with a single test regardless whether the alternative hypothesis is sparse or dense, Sardy et al. (2022) propose the $\oplus$-test.
The $\infty$-S test can be used to derive confidence regions that are dual to tests. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th canonical vector, and the calculated statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
The $\infty$-S test is available and the research is made reproducible in the Stest package in R, which can be downloaded from https://github.com/StatisticsL/Stest.
7 Acknowledgement
This work was partially supported by the Swiss National Science Foundation grant 200021E_213166 and the Academic Rising Star Project No. 2024-YJRC-LXY-005 of the National University of Defense and Technology, China.
Appendix A Proof of Theorem 1
Consider the primal cost function in (5). This primal cost being convex, the gap between the primal and dual cost functions is zero at optimality. The dual cost can be derived as follows
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}$
$=$ $\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p},\boldsymbol{u}\in{\mathbb{R}}^{n}}\|{\boldsymbol{u}}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\quad{\rm s.t.}\quad\boldsymbol{u}=\boldsymbol{y}-X\boldsymbol{\beta}$
$=$ $\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p},\boldsymbol{u}\in{\mathbb{R}}^{n}}\max_{{\boldsymbol{\omega}}}\|{\boldsymbol{u}}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}(\boldsymbol{u}-\boldsymbol{y}+X\boldsymbol{\beta})$
$=$ $\max_{{\boldsymbol{\omega}}}-{\boldsymbol{\omega}}^{\rm T}\boldsymbol{y}+\min_{\boldsymbol{u}\in{\mathbb{R}}^{n}}\|{\boldsymbol{u}}\|_{1}+{\boldsymbol{\omega}}^{\rm T}\boldsymbol{u}+\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}X\boldsymbol{\beta}$
$=$ $\max_{\|{\boldsymbol{\omega}}\|_{\infty}\leq 1,X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)}-{\boldsymbol{\omega}}^{\rm T}\boldsymbol{y}+0+\lambda\|A\boldsymbol{\beta}^{\star}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}X\boldsymbol{\beta}^{\star},$
where $\boldsymbol{\beta}^{\star}$ satisfies the KKT conditions
$\lambda A^{\rm T}\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})\ni X^{\rm T}{\boldsymbol{\omega}},$ (15)
where $\gamma()$ applied componentwise to the penalized vector ${\boldsymbol{\kappa}}=A\boldsymbol{\beta}^{\star}-\boldsymbol{b}$ is defined by $\gamma(\kappa_{k})={\rm sign}(\kappa_{k})$ if $\kappa_{k}\neq 0$, and $\gamma(0)\in[-1,1]$, for $k=1,\ldots,m$. We are interested in finding the smallest $\lambda$ for the solution to (5) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and (15) is
$\lambda A^{\rm T}{\cal B}^{m}_{\infty}(1)\ni X^{\rm T}{\boldsymbol{\omega}}.$ (16)
But $X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)$ and $A$ is full row rank, so there exists a unique ${\boldsymbol{\alpha}}\in{\mathbb{R}}^{m}$ such that $X^{\rm T}{\boldsymbol{\omega}}=A^{\rm T}{\boldsymbol{\alpha}}$; moreover ${\boldsymbol{\alpha}}=(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}$ is this unique solution. The smallest $\lambda$ satisfying conditions (16) is $\|{\boldsymbol{\alpha}}\|_{\infty}$, so
$\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}.$
Under $H_{0}$, the dual cost is $g({\boldsymbol{\omega}}^{\star}):=-{{\boldsymbol{\omega}}^{\star}}^{\rm T}\boldsymbol{y}+0+{{\boldsymbol{\omega}}^{\star}}^{\rm T}X\boldsymbol{\beta}^{\star}$, and the primal cost is $f(\boldsymbol{\beta}^{\star}):=\|\boldsymbol{y}-X\boldsymbol{\beta}^{\star}\|_{1}$ with $\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}_{H_{0}}$. Since the duality gap is zero, then $g({\boldsymbol{\omega}}^{\star})=f(\boldsymbol{\beta}^{\star})$, that is, $-{{\boldsymbol{\omega}}^{\star}}^{\rm T}(\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}})=\|\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}\|_{1}$. Consequently ${\boldsymbol{\omega}}^{\star}=\gamma(X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}})$, with the same definition of $\gamma()$ as above. Some entries {${{\cal I}\in\{1,\ldots,n\}}$ of the residuals ${\boldsymbol{r}}={\boldsymbol{y}}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ are different from zero, in which case ${\boldsymbol{\omega}}^{\star}_{\cal I}={\rm sign}(-{\boldsymbol{r}}_{I})$. Let now $K\in{\mathbb{R}}^{p\times(p-m)}$ be a basis for the kernel of $A$. We have that $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}={\boldsymbol{0}}_{p-m}$. Letting $X_{\cal I}$ be the rows of $X$ which indexes are in ${\cal I}$, $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}^{\star}=K^{\rm T}(X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}+X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}})\equiv 0$ iff $K^{\rm T}X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}=-K^{\rm T}X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}$. Solving this linear system leads to the remaining vector of the dual ${\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}\in{\mathbb{R}}^{p-m}$.
Appendix B Proof of Theorem 2
The residuals of the least absolute fit have a distribution which scale is proportional to $\xi$; so since $S=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the LAD estimate ${\hat{\boldsymbol{\beta}}_{H_{0}}}$ is asymptotically Gaussian with mean ${\boldsymbol{\beta}}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{\beta}_{H_{0}}})-X{\boldsymbol{\epsilon}}$ is asymptotically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}$.
Appendix C Proof of Theorem 4
It is well known that the LAD optimization problem (here with $\boldsymbol{y}=\boldsymbol{y}_{-A}$ and $X=X_{-A}$) can be rewritten as the linear program:
$\min_{\boldsymbol{r}_{+}\geq 0,\boldsymbol{r}_{-}\geq 0,\boldsymbol{\beta}}(\boldsymbol{r}_{+}^{\rm T},\boldsymbol{r}_{-}^{\rm T})^{\rm T}\boldsymbol{1}_{2n}\quad{\rm s.t.}\quad\boldsymbol{y}_{-A}=X_{-A}\boldsymbol{\beta}_{-A}+\boldsymbol{r}_{+}-\boldsymbol{r}_{-},$
where $\boldsymbol{1}_{2n}$ is the vector of ones of length $2n$. Using the Lagrange multiplier dual variable $\boldsymbol{\omega}$, the dual can be derived as follows.
$\min_{\boldsymbol{r}_{+}\geq 0,\boldsymbol{r}_{-}\geq 0,\boldsymbol{\beta}_{-A}}\max_{\boldsymbol{\omega}}(\boldsymbol{r}_{+}^{\rm T},\boldsymbol{r}_{-}^{\rm T})^{\rm T}\boldsymbol{1}_{2n}+\boldsymbol{\omega}^{\rm T}(\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{-A}-\boldsymbol{r}_{+}+\boldsymbol{r}_{-})$
$=$ $\max_{\boldsymbol{\omega}}\boldsymbol{\omega}^{\rm T}\boldsymbol{y}_{-A}-\min_{\boldsymbol{\beta}_{-A}}\boldsymbol{\beta}_{-A}^{\rm T}X_{-A}^{\rm T}\boldsymbol{\omega}+\min_{\boldsymbol{r}_{+}\geq 0}(\boldsymbol{1}_{n}^{\rm T}\boldsymbol{r}_{+}-\boldsymbol{\omega}^{\rm T}\boldsymbol{r}_{+})+\min_{\boldsymbol{r}_{-}\geq 0}(\boldsymbol{1}_{n}^{\rm T}\boldsymbol{r}_{-}+\boldsymbol{\omega}^{\rm T}\boldsymbol{r}_{-})$
$=$ $\max_{\|\boldsymbol{\omega}\|_{\infty}\leq 1,X_{-A}^{\rm T}\boldsymbol{\omega}={\bf 0}}\boldsymbol{\omega}^{\rm T}\boldsymbol{y}_{-A}+0$
with $w_{i}=1=$ when $(r_{+})_{i}>0$, $w_{i}=-1$ when $(r_{-})_{i}>0$, so $\boldsymbol{\omega}$ is the sign of the residuals $\boldsymbol{y}_{-A}-X_{-A}\hat{{\boldsymbol{\beta}}}_{-A}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ when the residuals are non-zero. So the dual problems of the constrained LAD and affine LAD LASSO when $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ are the same.
Appendix D Proof of Theorem 5
The proof follows the same lines as the proof of Theorem 1. The dual cost to (12) is the same except that the constraint on ${\boldsymbol{\omega}}\in[-\tau,1-\tau]^{n}$ with the same KKT conditions. We are interested in finding the smallest $\lambda$ for the solution to (12) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}^{\tau}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and the KKT conditions are
$\lambda A^{\rm T}{\cal B}^{m}_{\infty}(1)\ni X^{\rm T}{\boldsymbol{\omega}}.$ (17)
But $X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)$ and $A$ is full row rank, so there exists a unique ${\boldsymbol{\alpha}}\in{\mathbb{R}}^{m}$ such that $X^{\rm T}{\boldsymbol{\omega}}=A^{\rm T}{\boldsymbol{\alpha}}$; moreover ${\boldsymbol{\alpha}}=(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}$ is this unique solution. The smallest $\lambda$ satisfying conditions (17) is $\|{\boldsymbol{\alpha}}\|_{\infty}$, so
$\lambda_{0}^{\tau}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}.$
Under $H_{0}$, the dual cost is $g({\boldsymbol{\omega}}^{\star}):=-{{\boldsymbol{\omega}}^{\star}}^{\rm T}\boldsymbol{y}+0+{{\boldsymbol{\omega}}^{\star}}^{\rm T}X\boldsymbol{\beta}^{\star}$, and the primal cost is $f(\boldsymbol{\beta}^{\star}):=\|\boldsymbol{y}-X\boldsymbol{\beta}^{\star}\|_{\rho_{\tau}}$ with $\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}$. Since the duality gap is zero, then $g({\boldsymbol{\omega}}^{\star})=f(\boldsymbol{\beta}^{\star})$, that is, $-{{\boldsymbol{\omega}}^{\star}}^{\rm T}{\boldsymbol{r}}=\|\boldsymbol{r}\|_{\rho_{\tau}}$ with the residuals ${\boldsymbol{r}}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}$. Consequently ${\boldsymbol{\omega}}^{\star}=\gamma_{\tau}({\bf r})$ with $\gamma_{\tau}(r_{k})=1-\tau$ if $r_{k}<0$, $\gamma_{\tau}(r_{k})=-\tau$ if $r_{k}>0$ and $\gamma_{\tau}(0)\in[\tau-1,\tau]$, for $k=1,\ldots,m$. Some entries {${{\cal I}\in\{1,\ldots,n\}}$ of the residuals ${\boldsymbol{r}}$ are different from zero, in which case ${\boldsymbol{\omega}}^{\star}_{\cal I}=\gamma_{\tau}({\boldsymbol{r}}_{I})$. Let now $K\in{\mathbb{R}}^{p\times(p-m)}$ be a basis for the kernel of $A$. We have that $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}={\boldsymbol{0}}_{p-m}$. Letting $X_{\cal I}$ be the rows of $X$ which indexes are in ${\cal I}$, $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}^{\star}=K^{\rm T}(X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}+X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}})\equiv 0$ iff $K^{\rm T}X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}=-K^{\rm T}X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}$. Solving this linear system leads to the remaining vector of the dual ${\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}\in{\mathbb{R}}^{p-m}$.
Appendix E Proof of Theorem 6
The residuals of the least $\rho_{\tau}$ fit have a distribution which scale is proportional to $\xi$; so since $S^{\tau}=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the quantile regression estimate ${\hat{\boldsymbol{\beta}}^{\tau}_{H_{0}}}$ is asymptotically Gaussian with mean ${\boldsymbol{\beta}}^{\tau}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{\beta}_{H_{0}}^{\tau}})-X{\boldsymbol{\epsilon}}$ is asymptically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}^{\tau}$.