### v3

Abstract
A novel test in the linear $\ell_{1}$ (LAD) and quantile regressions is proposed, based on the scores provided by the dual variables (signs) arising in the calculation of the (so-called) affine-lasso estimate—a Rao-type, Lagrange multiplier test using the thresholding, towards the null hypothesis of the test, function of the latter estimate.
keywords:
LAD regression, quantile regression, LASSO, Nonparametric statistical test, Rescaling, Robustness, Total variation.
1 Introduction
We propose a new statistical test of a linear hypothesis in the linear $\ell_{1}$ (LAD) and quantile regression. In that context, the inference is considerably nonparametric, as the null hypothesis in testing is as a rule expressed via certain restriction on the median or pertinent quantile. For instance, the well-known sign test, going back to John Arbuthnot in 1710 (Spr89; Con99), tests the null hypothesis that the common median of independent observations $Y_{i}$ is $\mu$; the test statistic is then the number of positive signs among the signs of all $Y_{i}-\mu$ (in practice, one would like to avoid zeros of $Y_{i}-\mu$, but that is a minor complication that can be dealt with). Apart from the independence assumption, the distributional specification is quite loose; the $Y_{i}$’s even do not have to have the same distribution. Thus, there is a lot of peace of mind in the application of the test, as one has not to put too much faith in overly detailed assumptions. In the traditional terminology, the sign test is dubbed nonparametric, meaning that it is guaranteed to maintain the nominal level under fairly weak assumptions—and at the same time exhibits reasonable power against numerous alternatives; of course, here being nothing for free, the potential cost of such generality may be the loss of efficiency compared to certain parametric tests derived for exactly specified distributions—but only rather in case when the data really follow those.
It is well-known now that the use of the $\ell_{1}$ loss (cost) function in regression predated the introduction of the $\ell_{2}$ one by Gauss and Legendre by about half a century. The computation feasibility of the latter, combined with the flexibility and plausibility of the distributional assumption of normality, eclipsed the early 1760 invention of Boscovich for the centuries to come. However, the computational breakthroughs in linear programming, and among all, the possibility of broadening of the scope of the $\ell_{1}$ approach of the LAD (“Least Absolute Deviations”) regression to quantile regression, the study of conditional quantiles championed by KB78, brought the pendulum of the attention considerably back.
This gained an additional strong momentum when the $\ell_{1}$ loss function emerged also in the penalized formulations of LASSO (Least Absolute Shrinkage and Selection Operator, also known as Basis Pursuit)(CDS99; Tibs:regr:1996). While the original formulations retained the $\ell_{2}$ loss in the lack-of-fit term, to maintain a bond with the prevailing methodology, an attractive synthesis arose in the quantile regression LASSO, its particular median case known as LAD-LASSO (CIS-215377); the latter framework carries a conceptual advantage that both lack-of-fit term and penalty term feature the same form of the loss functions—so that the LAD-LASSO can be written in an augmented $\ell_{1}$ regression form.
All that said, the problem of inference in the $\ell_{1}$ regression is indisputably appealing. The large-sample approach dates back to kolmog31, who already established the asymptotic normality of the sample median for independent, identically distributed observations; this was extended to the regression case by LADasym91, and refined by others. The main hurdle of the implementation of this approach is that the asymptotic variance depends on the reciprocal of the density (that is, “sparsity”) at the (potentially unknown) median of the distribution of the observations (the distribution of the errors in the regression case). The large-sample Wald-type test is then bound to estimate this unknown nuisance parameter; while this is not an impossible task—various strategies proposed in this respect are reviewed in the book of Koenker—it may still constitute a hurdle one would like to bypass.
And indeed: the Rao’s “score” approach to testing (also known as Lagrange multiplier test), elaborated for inference in the $\ell_{1}$/quantile regression by Roger93 via so-called regression rank scores (gutejure92), the dual variables of the primal convex optimization task defining the estimates, brought far-reaching generalizations of the traditional one-sample and two-sample rank tests (including, in particular, the sign test). While it is a general wisdom that these methods avoid the estimation of sparsity—as underlined explicitly by Koenkerin the context of confidence intervals—he2023 mention computational instabilities for large samples and various other problems, and also the estimation of asymptotic variance which could allegedly mar the rank test approach—quoting, however, rather the more abstract account of gutejure92 instead of the more focused Roger93 .
The problem here is pretty much in the eye of beholder, however. Unlike the simple hit-and-miss paradigm of the sign-test, rank tests allow for various flavors depending on what kind of rank scores are adopted; the scores make Wilcoxon tests different from the van der Waerden ones, for example. A human investigator may prefer one to another for the one- and two-sample problems, and accordingly may opt for the corresponding scores also in the $\ell_{1}$/quantile setting. For instance, van der Waerden scores may be preferred by those who believe that the reality is mostly normal (“Gaussian”), apart from occasional erratic disturbances. Thus, once the flavor of the rank test is chosen a priori, the procedure “does not require any nuisance parameter depending on the error distribution to be estimated” (Roger93).
However, when the rank testing is to be performed in an automatic way, the rank scores may have to be elucidated in an “adaptive way”, using the existing theory of optimality of certain scores for certain distributions (for example, as mentioned above, van der Waerden scores work best for the normal distribution) to estimate the right scores directly from the data. Ascending this next step on the ladder brings a need of distributional estimation, potentially including that of sparsity, back into the game.
The desire for automatic, human-free procedures may thus favor a simple solution without a need of additional tuning choices; if such a solution is available, and exhibits decent power compared to the already existing options, it possesses a definite raison-d’être, the aspect we aim at demonstrating below. Section 2 gives the derivation of $\infty$-S test by defining the affine LAD LASSO, elucidating its zero-thresholding function, and subsequently proposing the sign-based test statistic that is asymptotically pivotal with respect to the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO; Section 4 extends the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile affine LASSO estimator. In Section 5, the $\infty$-S test is applied in four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
2 The $\infty$-S test
Consider a linear model
$\boldsymbol{y}=X\boldsymbol{\beta}+\boldsymbol{e},$ (1)
where $X$ is an $n\times p$ matrix and $\boldsymbol{\beta}$ is a $p\times 1$ vector of unknown parameters. Given an $m\times p$ matrix $A$ of rank $m$ and a vector $\boldsymbol{b}\in{\mathbb{R}}^{m}$, we consider testing
$H_{0}:\ A\boldsymbol{\beta}=\boldsymbol{b}\quad\text{against}\quad H_{1}:\ A\boldsymbol{\beta}\neq\boldsymbol{b}.$ (2)
A new testing procedure, which we propose to call an $\infty$-S test, is based on the least absolute deviation version of the affine LASSO point estimator (SARDY2022107507), a regularized estimator that can be defined in the constrained form as a solution of the problem
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!\qquad\text{subject to}\qquad\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\leq\Lambda.$ (3)
The “Lagrangian” form of the definition used by SARDY2022107507 proclaims the same estimator a solution to
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (4)
for a suitable Lagrange multiplier $\lambda$ pertaining to the constraint in (3); the larger $\lambda$, the more this estimator thresholds towards zero the entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ corresponding to the null hypothesis $H_{0}$ in (2). That is, there is an interval $\lambda_{0},\infty)$ for which the solution to (4) satisfies the condition of $H_{0}$ for any $\lambda\in[\lambda_{0},\infty)$. The smallest such $\lambda$ is a function of the data and has a closed form expression which we call the zero-thresholding function.
Theorem 1.
Let
$\hat{\boldsymbol{\beta}}_{H_{0}}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\quad\text{\rm subject to}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (5)
The zero-thresholding function of the affine LAD-LASSO primal problem (4) is
$\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty},$ (6)
where $\boldsymbol{\omega}$} is the dual variable associated to (4), the sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}$ of the constrained LAD (5) whenever $X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}}\neq 0$.
As satisfying $H_{0}$ is equivalent to $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$ for any norm, the result of the zero-thresholding function (6) is the the constraint $\|A\boldsymbol{\beta}-\boldsymbol{b}\|=0$; with the $\ell_{1}$-norm, the pertinent Lagrange multiplier is finite. Instead of the $\ell_{1}$-norm, we could have used the $\ell_{2}$-norm corresponding to affine LAD-group LASSO (Yuan:Lin:mode:2006); the dual of the $\ell_{2}$-norm being the $\ell_{2}$-norm itself, the corresponding zero-thresholding function would have been $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{2}$. More generally, one can use the $\ell_{q}$-norm, which corresponding zero-thresholding function is $\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{q/(q-1)}$.
The zero-thresholding function yields the test statistic of the $\infty$-S test, which the following theorem shows is asymptotically pivotal. The $\infty$-S test can be thus considered “Lagrange multiplier test”—that is, the Rao score test— akin to the rank tests developed by Roger93; but, unlike the statistic $T_{n}$ of Roger93, the $\infty$-S test does not require the integration of scores for dual variables coming from different quantile regressions, and avoids also the need to select the rank scores (and subsequent potential estimation of underlying density characterizations to achieve the optimality of those).
Theorem 2.
Let $\boldsymbol{Y}$ be the response random vector under $H_{0}$. Assuming the LAD coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ is asymptotically pivotal.
The asymptotic normality of the LAD coefficient estimates required by Theorem 2 is a well known fact—proved by LADasym91, and refined by others. We are therefore in the position to define the $\infty$-S test now.
Definition 1 (The $\infty$-S test).
The $\infty$-S test function, to test (2) in a linear model  (1) at a prescribed level $\alpha\in(0,1)$, is defined to be
$\phi({\boldsymbol{y}},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}(\boldsymbol{y},X,A)\leq c_{\alpha},\\ 1&\text{\rm otherwise},\end{array}\right$.,
where $c_{\alpha}$ is selected so that the test has nominal level $\alpha$.
Based on Theorem 2, the level of the $\infty$-S test matches asymptotically the nominal level $\alpha$. The test itself is being based on the dual variables for the $\ell_{1}$ minimization problem; these variables attain values in [1,1][-1,1] and amount to the sign of residuals if those are not equal to zero. The insensitivity to the magnitude of the residuals makes the test resistant to outlying observations; nonetheless, in model-behaved situations, when the error distribution is exactly normal, a more powerful $\infty$-rankS test can be obtained by weighting the sign $\boldsymbol{\omega}$ of the residuals by the rank of the absolute value of the respective residuals.
In the applications, the sought p-values (or critical values $c_{\alpha}$, if desired) are then easily obtained via Monte Carlo sampling of $S=\lambda_{0}(\boldsymbol{Y},X,A)$, the technology which nowadays provides satisfactory results even in the absence (so far) of a knowledge of the distribution of test statistics in closed form (related, say, to some traditional distributions).
From the practical point of view, we also recommend, in the spirit of SardyISI08, to perform preliminary rescaling of $(A,\boldsymbol{b})$, in order to achieve homogeneous power—that is, the power not favoring certain alternative hypotheses to the detriment of the other ones. Indeed, the matrices $A$ and $X$ create heteroskedasticity in the random vector $\boldsymbol{W}$ in the test statistics $S=\|\boldsymbol{W}\|_{\infty}$ with $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ before taking the maximum of all entries. Consequently, under $H_{1}$, the components of $\boldsymbol{W}$ with small deviations will be hidden behind the components with large deviations. To treat equally potential components of $A\boldsymbol{\beta}-\boldsymbol{b}$ not being equal to zero, we propose the following.
Definition 2 (Homopower rescaling).
Consider the $\infty$-S test with test statistic $S=\|\boldsymbol{W}\|_{\infty}$, where $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ and $\boldsymbol{\omega}$} are the sign of the residuals of the constrained LAD under $H_{0}$. Let $F_{W_{k}}$ be the marginal cumulative distribution function of the $k$th component of $\boldsymbol{W}$ for $k=1,\ldots,m$. Letting $D_{(A,X,{\bf b},\alpha)}$ be the diagonal matrix with $k$th diagonal element equal to $d_{k}:=F_{W_{k}}^{-1}(1-\alpha)$, the homopower rescaling rescales $A$ to $D_{(X,A,{\bf b},\alpha)}A$ and $\boldsymbol{b}$ to $D_{(X,A,{\bf b},\alpha)}\boldsymbol{b}$ before applying the test.
3 Optimization aspects
The objective function of the LAD optimization task
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (7)
is convex, and it is well known that the optimization can be transformed into a linear program (KB78). The $\infty$-S test requires solving the constrained LAD optimization (5) that can be transformed into an unconstrained LAD optimization of the form (7).
Theorem 3.
Solving (5) is equivalent to solving (7) with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$, where $A_{1}$ are $m$ linearly independent columns of $A=:[A_{1},A_{2}]$ and $X=:[X_{1},X_{2}]$ is the corresponding decomposition.
The proof is straightforward: choose $m$ linearly independent columns of $A$, call them $A_{1}$ and the remaining ones $A_{2}$, so that $A=[A_{1},A_{2}]$, permuting $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ accordingly. The linear constraints thus lead to $\boldsymbol{\beta}_{1}=A_{1}^{-1}(\boldsymbol{b}-A_{2}\boldsymbol{\beta}_{2})$ and $\boldsymbol{y}-X\boldsymbol{\beta}$ becomes $\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{2}$ with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$.
The affine LAD-LASSO optimization (4) can be also transformed into an unpenalized LAD of the form (3), as it is equivalent to solving
$\min_{{\boldsymbol{\beta}}\in{\mathbb{R}}^{p}}\biggl\|\left(\begin{array}[]{c}{\boldsymbol{y}}\\ {\lambda{\boldsymbol{b}}}\end{array}\right)-\left(\begin{array}[]{c}X\\ {\lambda A}\end{array}\right){\boldsymbol{\beta}}\biggr\|_{1}.$
Another way to perform the constrained LAD optimization (5) is to solve the affine LAD-LASSO (4) for $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ of Theorem 1 to guarantee complete thresholding—that is, enforcing the linear constraint; this is akin to the exact penalty method of PG89. The following theorem links the dual of the affine LAD-LASSO optimization (4) and the constrained LAD optimization.
Theorem 4.
The dual variable $\boldsymbol{\omega}$ of Theorem 1 corresponding to the primal affine LAD LASSO optimization (4) is the dual variable of the LAD optimization with $\boldsymbol{y}_{-A}$ and $X_{-A}$ of Theorem 3.
Summing up, the $\infty$-S test can be straightforwardly implemented using the existing software—in particular, employing the R package quantreg (Rogerrq).
4 The $\infty$-Sτ test for quantile regression
Given $\tau\in(0,1)$, let ${\boldsymbol{q}}^{\tau}$ be conditional $\tau$-quantile vector ${\boldsymbol{q}}^{\tau}$ of a response vector ${\boldsymbol{Y}}\mid X$. The linear model assumes the form
${\boldsymbol{q}}^{\tau}=X{\boldsymbol{\beta}},$ (8)
where $X$ is the regression matrix as before. To estimate the parameters $\boldsymbol{\beta}$} from $\boldsymbol{y}$} arising from the response random vector $\boldsymbol{Y}$}, quantile regression (KB78) solves the optimization problem
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{\rho_{\tau}}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (9)
The residuals $\boldsymbol{r}=\boldsymbol{y}-X\boldsymbol{\beta}$ are now subjected to the objective function $\|\boldsymbol{r}\|_{\rho_{\tau}}=\sum_{i=1}^{n}\rho_{\tau}(r_{i})$, where
$\rho_{\tau}(r_{i})=\left\{\begin{array}[]{rl}(\tau-1)r_{i}&\ r_{i}<0,\\ \tau r_{i}&\ r_{i}\geq 0,\end{array}\right$. (10)
is the “check function”—the tilted $\ell_{1}$, “pinball” loss. Note that the special case of $\tau=1/2$ reduces to the LAD regression, as $\|\boldsymbol{r}\|_{\rho_{1/2}}=\|{\boldsymbol{r}}\|_{1}/2$.
To test the null hypothesis against the alternative hypothesis, as in (2) but in the quantile-regression spirit, the quantile $\infty$-S test does not refer to the median (well, unless $\tau=1/2$), but rather to (any) $\tau$-quantile for some fixed $\tau\in(0,1)$. We define accordingly the affine $\rho_{\tau}$-LASSO as a solution to
$\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (11)
for fixed $\lambda>0$ and $\tau\in(0,1)$. The following theorem gives the zero-thresholding function of this estimator.
Theorem 5.
For fixed $\tau\in(0,1)$, let
$\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})\quad{\rm subject\ to}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (12)
The zero-thresholding function of the affine $\rho_{\tau}$-LASSO primal problem (11) is
$\lambda_{0}^{\tau}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty},$ (13)
where ${\boldsymbol{\omega}}^{\tau}$ is the dual variable associated to the linear programming solution of (11).
The zero-thresholding function leads to the statistic of the $\infty$-Sτ test; the following theorem shows that it is asymptotically pivotal.
Theorem 6.
Let $\boldsymbol{Y}$ be the response random vector under $H_{0}$ and let $\tau\in(0,1)$. If the $\tau$-quantile regression parameter estimates are asymptotically normal with the limit distribution centered about their true values, the test statistic $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ is asymptotically pivotal.
The asymptotic normality of the quantile regression estimators, as established by KB78 (their results including the $\ell_{1}$ results as a special case), allows for the following definition of the $\infty$-Sτ test—which Theorem 6 shows has asymptotically the nominal level $\alpha$.
Definition 3 (The $\infty$-Sτ test).
Given $\tau\in(0,1)$, the $\infty$-Sτ test to test (2) in the linear model (8) at a prescribed level $\alpha\in(0,1)$, is defined to be
$\phi({\boldsymbol{y}},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}^{\tau}(\boldsymbol{y},X,A)\leq c_{\alpha}^{\tau}\\ 1&{\rm otherwise}\end{array}\right$.,
where $c^{\tau}_{\alpha}$ is selected so that the test has nominal level $\alpha$.
5 Special cases and applications
Our simulations and applications rely on the R package quantreg (Rogerrq). We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ involved in the zero-thresholding function. In all our applications of the $\infty$-S test, the critical values are estimated by Monte Carlo simulation with $10^{4}$ runs. The levels and power curves of the tests are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the $\chi^{2}$ sign test whenever it returns a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
5.1 The paired $\infty$-S test
Consider two paired measurements $\{(u_{i},v_{i})\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu_{i}+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu_{i}+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
where the errors $\tilde{\epsilon}$ and $\breve{\epsilon}$ are assumed independent with a median equal to zero. To test
$H_{0}:\delta=0\quad{\rm against}\quad H_{1}:\delta\neq 0$
with the $\infty$-S test, we write the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{v}}-{\boldsymbol{u}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. The value of the test statistic is then
$s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{v}}-{\boldsymbol{u}})|=|\sum_{i=1}^{n}(1(v_{i}>u_{i})-1(v_{i}<u_{i}))|.$
Given that the statistic of the sign test statistic in this situation is
$\tilde{s}=\sum_{i=1}^{n}1(v_{i}>u_{i})\quad\text{and}\quad\sum_{i=1}^{n}1(v_{i}>u_{i})+\sum_{i=1}^{n}1(v_{i}<u_{i})=n,$
we have that $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the sign test.
5.2 The unpaired $\infty$-S test
Consider two unpaired measurements $\{u_{i}\}_{i=1,\ldots,n}$ and $\{v_{i}\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
where the errors $\tilde{\epsilon}$ and $\breve{\epsilon}$ are assumed independent with a median equal to zero. To test
$H_{0}:\delta=0\quad{\rm against}\quad H_{1}:\delta\neq 0$
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X{\boldsymbol{\beta}}+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}^{\rm T}=({\boldsymbol{u}}^{\rm T},{\boldsymbol{v}}^{\rm T})$,
$X=\left(\begin{array}[]{cc}\begin{array}[]{c}{\boldsymbol{1}}_{n}\\ {\boldsymbol{1}}_{n}\end{array}&\begin{array}[]{c}{\boldsymbol{0}}_{n}\\ {\boldsymbol{1}}_{n}\end{array}\end{array}\right),$
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets now $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is
$s=\max(0,|\sum_{i=1}^{n}1(v_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(v_{i}<m_{\boldsymbol{y}})|).$
We obtain the sign test for unpaired data, also known as the median test (brown1951median).
Figure 1: Power of the unpaired $\infty$-Sτ test (same as median test) and $\infty$-rankSτ test for $\tau=0.5$, the F-test and asymptotically $\chi^{2}$ sign test as a function of the shift $\delta\in[0,2]$ between the two populations. The horizontal dotted line is the nominal level $\alpha=0.05$. Sample size $n=20$ (left plot) and $n=100$ (right plot).
We compare the level and the power of four tests ($\infty$-S or median test, $\infty$-rankS test, F-test and $\chi^{2}$ sign test) on two Monte Carlo simulations with $n=20$ and $n=100$ and with Student errors with 3 degrees of freedom. Figure 1 shows that all tests satisfy the nominal level of $\alpha=0.05$ and the F-test looks quite robust to non-Gaussian errors although losing some power. The $\chi^{2}$ sign test seems to overshoot the level, even for $n$ large, however.
5.3 Homopower rescaling
We perform a small simulation study to outline the importance of homopower rescaling (see Definition 2) of the two components $A$ and $\boldsymbol{b}$ of the linear null hypothesis (2). We set $n=100$, $p=20$, $m=2$, $\tau=0.5$, $\boldsymbol{b}=\boldsymbol{0}_{m}$ and $A=[C,O_{m\times(p-m)}]$ with
$C=\left(\begin{array}[]{ccc}3&0\\ 0&1\end{array}\right).$
The regression matrix $X$ has entries sampled from a standard normal random distribution, the null hypothesis is $H_{0}:\ \boldsymbol{\beta}=\boldsymbol{0}$; we consider two alternative hypotheses
$H_{1,1}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{1}\quad{\rm and}\quad H_{1,2}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{2},$
where $\boldsymbol{e}_{i}$ is the $i$th coordinate (canonical) vector. With the unrescaled $\infty$-S test, one expects the power to be low under $H_{1,1}$ and good under $H_{1,2}$, while with the rescaled $\infty$-S test, one expects the power to be high under both $H_{1,1}$ and $H_{1,2}$. Figure 2 illustrates that it is really so. While this example represents certainly just one very special case, we remark that what we observe here also occurs naturally, owing to the relation between $A$ and $X$ that is not controlled by the user (except possibly in ANOVA). Thus, it is not only the $\infty$-S test, but many other tests that would profit from such a rescaling in terms of power.
Figure 2: Homopower rescaling. Top: boxplots of Monte Carlo simulated bivariate ($m=2$) random vector $W$ unrescaled (left) and rescaled (right). Bottom: power plots of the unrescaled (red) and rescaled (black) $\infty$-S test under $H_{1,1}$ (left) and $H_{1,2}$ (right).
5.4 The total variation $\infty$-Sτtest
Total variation (ROF92) aims at detecting jumps in a noisy signal or a time series ${\bf y}=\boldsymbol{\beta}+\boldsymbol{e}$, which corresponds to model (1) with $X=I_{n}$. A jump occurs if $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\neq 0$. Let $A$ be the $(n-1)\times n$ (sparse) matrix with 1-1 and $1$ entries such that
$\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\ =\|A\boldsymbol{\beta}\|_{1}.$
Note that $A$ is of full row rank. To test for jump in the median or any $\tau$-quantile, one can apply the $\infty$-Sτ test or the $\infty$-rankSτ test at the desired $\tau$-quantile.
We apply the quantile total variation $\infty$-Sτ test to analyze the time series of the Amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of $0.0132$. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic (by Monte Carlo sampling of the test statistic $S^{\tau}$ under the null hypothesis) and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Right below this plot, Figure 3 shows the quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the quantile affine LASSO estimate for $\lambda=c_{\alpha}^{\tau}$ (bottom), which corresponds to the quantile universal threshold estimate (Giacoetal17) of the $\tau$-quantile. The single jump points to a potential change of regime moving to a less volatile return.
Figure 3: AMZN daily log-returm time series analysis. Top left plot: estimated density function of $f_{S^{\tau}}$ and critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$ by Monte Carlo simulation. Bottom left: regression quantile affine LASSO path for $\tau=0.1$ with critical value $c_{\alpha}^{\tau}$. Top right plot: AMZN time series data and estimate jump locations. Bottom right: estimated $\tau$-quantiles.
5.5 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, AA is the first $m=5$ rows of the finite difference matrix of Section 5.4 and $\boldsymbol{\beta}$} are $p$ samples from the standard Gaussian distribution; $\boldsymbol{b}$ is calculated as $A{\boldsymbol{\beta}}$. In a robustness analysis of the F-test, ALI96 observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom.
Figure 4: Level of the $\infty$-S test, $\infty$-rankS test and F-test as a function of the degrees of freedom of the Student errors. The dotted line is the nominal level $\alpha=0.01$.
Over-shooting the nominal level leads to too many type I errors, hence rejecting too many null hypotheses, that is, making too many false discoveries. Here the two $\infty$-S tests match the nominal level, regardless of the degree of freedom of the Student errors. For this simulation, the implementation of the quantreg package does not return a p-value for $\chi^{2}$ sign test.
6 Conclusions
The $\infty$-Sτ test allows to test general linear null hypotheses for linear models and for any quantile $\tau$. The test function is simple to implement and can handle high-dimensional problems. We emphasize the $\infty$-S test which corresponds to the $\ell_{1}$-LASSO penalty, and showed that other norms than the $\ell_{1}$-norm could be employed, leading to other tests. SARDY2022107507 show that using $\ell_{1}$ leads to high power under sparse alternatives (that is, most entries, but not all, of $A\boldsymbol{\beta}-\boldsymbol{b}$ are zero when $H_{1}$ is true), and, under dense alternatives (that is, most or all entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ are non-zero), that the sign test based on $\ell_{2}$ has more power than the $\infty$-S test. To get nearly the best power with a single test regardless whether the alternative hypothesis is sparse or dense, SARDY2022107507 propose the $\oplus$-test.
The inversion of $\infty$-S test can be used to derive confidence regions. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th coordinate (canonical) vector; the statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
The $\infty$-S test is available and the research is made reproducible in the Stest package in R, which can be downloaded from https://github.com/StatisticsL/Stest.
[…]
Acknowledgements
The research of Sylvain Sardy was partially supported by the Swiss National Science Foundation grant 200021E_213166, a partner in the Lead Agency project GAČR No. 23–06461K of the Czech Science Foundation, which supported the research of Ivan Mizera; the research of Xiaoyu Ma was supported by the National Natural Science Foundation of China(No. 12401397).
Appendix A Proof of Theorem 1
Consider the primal cost function in (2.5). This primal cost being convex, the gap between the primal and dual cost functions is zero at optimality. The dual cost can be derived as follows
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}$
$=$ $\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p},\boldsymbol{u}\in{\mathbb{R}}^{n}}\|{\boldsymbol{u}}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\quad{\rm s.t.}\quad\boldsymbol{u}=\boldsymbol{y}-X\boldsymbol{\beta}$
$=$ $\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p},\boldsymbol{u}\in{\mathbb{R}}^{n}}\max_{{\boldsymbol{\omega}}}\|{\boldsymbol{u}}\|_{1}+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}(\boldsymbol{u}-\boldsymbol{y}+X\boldsymbol{\beta})$
$=$ $\max_{{\boldsymbol{\omega}}}-{\boldsymbol{\omega}}^{\rm T}\boldsymbol{y}+\min_{\boldsymbol{u}\in{\mathbb{R}}^{n}}\|{\boldsymbol{u}}\|_{1}+{\boldsymbol{\omega}}^{\rm T}\boldsymbol{u}+\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}X\boldsymbol{\beta}$
$=$ $\max_{\|{\boldsymbol{\omega}}\|_{\infty}\leq 1,X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)}-{\boldsymbol{\omega}}^{\rm T}\boldsymbol{y}+0+\lambda\|A\boldsymbol{\beta}^{\star}-\boldsymbol{b}\|_{1}+{\boldsymbol{\omega}}^{\rm T}X\boldsymbol{\beta}^{\star},$
where $\boldsymbol{\beta}^{\star}$ satisfies the KKT conditions
$\lambda A^{\rm T}\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})\ni X^{\rm T}{\boldsymbol{\omega}},$ (14)
where $\gamma()$ applied componentwise to the penalized vector ${\boldsymbol{\kappa}}=A\boldsymbol{\beta}^{\star}-\boldsymbol{b}$ is defined by $\gamma(\kappa_{k})={\rm sign}(\kappa_{k})$ if $\kappa_{k}\neq 0$, and $\gamma(0)\in[-1,1]$, for $k=1,\ldots,m$. We are interested in finding the smallest $\lambda$ for the solution to (2.5) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and (14) is
$\lambda A^{\rm T}{\cal B}^{m}_{\infty}(1)\ni X^{\rm T}{\boldsymbol{\omega}}.$ (15)
[…]
Appendix B Proof of Theorem 2
The residuals of the least absolute fit have a distribution which scale is proportional to $\xi$; so since $S=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the LAD estimate $\hat{\boldsymbol{\beta}}_{H_{0}}$} is asymptotically Gaussian with mean ${\boldsymbol{\beta}}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{\beta}_{H_{0}}})-X{\boldsymbol{\epsilon}}$ is asymptotically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}$.
Appendix C Proof of Theorem 4
It is well known that the LAD optimization problem (here with $\boldsymbol{y}=\boldsymbol{y}_{-A}$ and $X=X_{-A}$) can be rewritten as the linear program: min_r_+≥0, r_-≥0,β (r_+^T, r_-^T)^T 1_2n s.t. y_-A=X_-Aβ_-A+r_+-r_-, where $\boldsymbol{1}_{2n}$ is the vector of ones of length $2n$. Using the Lagrange multiplier dual variable $\boldsymbol{\omega}$, the dual can be derived as follows.
$\min_{\boldsymbol{r}_{+}\geq 0,\boldsymbol{r}_{-}\geq 0,\boldsymbol{\beta}_{-A}}\max_{\boldsymbol{\omega}}(\boldsymbol{r}_{+}^{\rm T},\boldsymbol{r}_{-}^{\rm T})^{\rm T}\boldsymbol{1}_{2n}+\boldsymbol{\omega}^{\rm T}(\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{-A}-\boldsymbol{r}_{+}+\boldsymbol{r}_{-})$
$=$ $\max_{\boldsymbol{\omega}}\boldsymbol{\omega}^{\rm T}\boldsymbol{y}_{-A}-\min_{\boldsymbol{\beta}_{-A}}\boldsymbol{\beta}_{-A}^{\rm T}X_{-A}^{\rm T}\boldsymbol{\omega}+\min_{\boldsymbol{r}_{+}\geq 0}(\boldsymbol{1}_{n}^{\rm T}\boldsymbol{r}_{+}-\boldsymbol{\omega}^{\rm T}\boldsymbol{r}_{+})+\min_{\boldsymbol{r}_{-}\geq 0}(\boldsymbol{1}_{n}^{\rm T}\boldsymbol{r}_{-}+\boldsymbol{\omega}^{\rm T}\boldsymbol{r}_{-})$
$=$ $\max_{\|\boldsymbol{\omega}\|_{\infty}\leq 1,X_{-A}^{\rm T}\boldsymbol{\omega}={\bf 0}}\boldsymbol{\omega}^{\rm T}\boldsymbol{y}_{-A}+0$
with $w_{i}=1=$ when $(r_{+})_{i}>0$, $w_{i}=-1$ when $(r_{-})_{i}>0$, so $\boldsymbol{\omega}$ is the sign of the residuals $\boldsymbol{y}_{-A}-X_{-A}\hat{{\boldsymbol{\beta}}}_{-A}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ when the residuals are non-zero. So the dual problems of the constrained LAD and affine LAD-LASSO when $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ are the same.
Appendix D Proof of Theorem 5
The proof follows the same lines as the proof of Theorem 1. The dual cost to (4.12) is the same except that the constraint on ${\boldsymbol{\omega}}\in[-\tau,1-\tau]^{n}$ with the same KKT conditions. We are interested in finding the smallest $\lambda$ for the solution to (4.12) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}^{\tau}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and the KKT conditions are
$\lambda A^{\rm T}{\cal B}^{m}_{\infty}(1)\ni X^{\rm T}{\boldsymbol{\omega}}.$ (16)
[…]
Appendix E Proof of Theorem 6
The residuals of the least $\rho_{\tau}$ fit have a distribution which scale is proportional to $\xi$; so since $S^{\tau}=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the quantile regression estimate $\hat{\boldsymbol{\beta}}^{\tau}_{H_{0}}$} is asymptotically Gaussian with mean ${\boldsymbol{\beta}}^{\tau}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{\beta}_{H_{0}}^{\tau}})-X{\boldsymbol{\epsilon}}$ is asymptotically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}^{\tau}$.
Sylvian Sardy, Section of Mathematics, University of Geneva, rue du Conseil-Général 7-9, 1205 Geneva, Switzerland.
Ivan Mizera, Department of Probability and Mathematical Statistics, Faculty of Mathematics and Physics, Charles University Prague, Sokolovská 83, 186 75 Praha, Czechia.
Xiaoyu Ma, College of Science, National University of Defense and Technology, 1 Fuyuan Road, 410000 Changsha, China.
Hugo Gaible, Department of Mathematics, École Normale Supérieur Paris-Saclay, 4 avenue des Sciences, 91190 Gif-sur-Yvette, France.